# Error Mitigation Thresholds in Noisy Random Quantum Circuits

### v4

@@ I Introduction
Based on these connections to disordered Ising models, our central prediction is that PEC and TEM have an error mitigation threshold in random circuits in spatial dimensions $D\geq 2$ when the noise is imperfectly characterized, i.e., there is uncertainty in the parameters of the noise model that have random fluctuations in space and/or time. We consider a model of random unitary circuits subject to depolarizing noise with binary disorder in the depolarization rate [see Fig. 1(a)], building on extensive work characterizing quantum phases and phase transitions in hybrid random circuits Fisher et al. (2022); Potter and Vasseur (2022). To mitigate the effects of the noise we apply a uniform “antinoise” channel that inverts the depolarizing noise on average. We show that a replica statistical mechanics description of the problem has close parallels to the classical random field Ising model (RFIM) in $D+1$ dimensions. At the zero-mean field condition in the RFIM, a simple heuristic argument, originally due to Imry and Ma, Imry and Ma (1975) shows that the ordered phase survives random symmetry-breaking terms in sufficiently high dimensions. In the random circuit problem, this analysis indicates the possibility for thresholds in the robustness of error mitigation for $D\geq 2$, while $D=1$ is the marginal dimension. We also analyze the performance of error mitigation as a function of circuit depth. For depths larger than the linear size of the system, the mitigation fails, resulting in a fidelity worse then one obtains for the maximally mixed state. The corresponding phase diagram of the system is shown in Fig. 1(b) for $D\geq 2$. Analytic and numerical studies of the two-copy replica theory and exact simulations of the mitigated circuit dynamics are consistent with our predictions.
@@
II Results
II.1 Setup
@@
where $\operatorname{tr}_{\bm{x}}[\rho]$ is a partial trace operation at site $\bm{x}$. Tuning $q_{a}=q_{\bm{x},t}$ inverts the noise at space-time site $(\bm{x},t)$ such that $\mathcal{A}_{\bm{x}}\circ\mathcal{E}_{\bm{x},t}(\rho)=\rho$. To include the disorder, the noise rates are drawn randomly at each space-time location of the circuit from one of two values $q_{\bm{x},t}=q_{1(2)}$ with probability $p$ and $1-p$, respectively, and $q_{1}<q_{2}$. The antinoise can be tuned to a zero mean-field condition by taking $(1-q_{a})=(1-q_{1})^{p}(1-q_{2})^{1-p}$ equal to a geometric mean. This condition ensures that a Pauli operator $\sigma_{\bm{x}}^{\mu}$ ($\mu\in\{X,Y,Z\}$) evolved under multiple rounds of noise and antinoise up to depth $t=d$ will have expectation value
@@
II.2 Mean-field theory
@@
We focus on the disordered case, however, where the noise rates are not known perfectly to the antinoise. In particular, $\gamma_{i}=\gamma_{1(2)}$ is drawn randomly with probability $p$ and $1-p$, respectively. The noise is canceled on average by setting $\gamma_{a}=\bar{\gamma}\equiv p\,\gamma_{1}+(1-p)\,\gamma_{2}$. We take the unitary evolution to be given by an all-to-all coupled Brownian circuit model in which the $J_{ij\mu\nu}(t)$ are drawn from a white-noise Gaussian random process with variance parameters Lashkari et al. (2013); Jian et al. (2022)
@@
To treat this model analytically, we average the dynamics over the random coupling constants for multiple identical copies of the density matrix $\rho(t)^{\otimes k}$ ($k$ is a replica index). In the equations of motion for the averaged moments $\rho_{k}(t)\equiv\mathbb{E}_{H}[\rho^{\otimes k}]$, the unitary drive-term averages to zero, leading to a purely dissipative master equation for $\rho_{k}(t)$ (see Methods A.1). Focusing on second moments $(k=2)$ gives rise to a particularly simple model with two global product steady states given by $I^{\otimes N}$ and $S^{\otimes N}$ Jian et al. (2022), where $S$ is the SWAP operator acting on the two copies. The all SWAP steady-state captures non-trivial corrections to second moment observables like the purity of sub-systems. In the mean-field approximation, we enforce the density matrix to take a product state form $\rho_{2}=\bigotimes_{i=1}^{N}\sigma_{i}$. Due to the lack of symmetries or conservation laws in the problem, $\sigma_{i}$ can be parameterized as
@@
where $\Delta_{i}=\gamma_{i}-\gamma_{a}$. This simple equation captures much of the basic physics we find later in the random circuit models, including the presence of a phase transition. Notably, we can see that at weak values of the disorder, the interaction term acts as a restoring force towards the fixed point $\delta_{i}=0$. The other fixed point $\delta_{i}=-3/4$ corresponds to an unphysical density matrix, which is unstable at weak disorder. At a critical disorder strength of $|\Delta_{1}|=3J$ (see App. A.1), however, the $\delta_{i}=0$ fixed point becomes unstable. The dynamics flows to a new stable fixed point which is an unphysical density matrix. Thus, we see that the above threshold phase in the mean-field theory is characterized by an instabilty of the density matrix to unphysical states.
@@
II.3 Statistical Mechanics Mapping
To analyze the problem more rigorously, we turn to well-established mappings from random quantum circuit dynamics in $D$ spatial dimensions to statistical mechanics models in $D+1$ dimensions Nahum et al. (2018); von Keyserlingk et al. (2018); Jian et al. (2020); Bao et al. (2020); Dalzell et al. (2022, 2021); Deshpande et al. (2022); Gao et al. (2021). We outline the details of the mapping in the Methods A.2. The approach we take is to expand $\rho_{k}\equiv\mathbb{U}[\rho^{\otimes k}]$ into a basis of operators based on representations of the permutation group $S_{k}$ Dalzell et al. (2022). One can then derive update rules for the averaged state following each layer of gates, noise, and antinoise. In this paper, we focus on the two-replica case where the two operators are $I$ and $S$ (the SWAP gate). We expect this model to be sufficient to capture many of the properties of the transition, as has been argued in the unitary case Nahum et al. (2018). Numerical evidence based on exact simulations is presented below that support this claim.
@@ II.3 Statistical Mechanics Mapping
Figure 2: Numerical tests of the threshold behavior: (a) Numerical simulations of circuit and noise-disorder averaged the correlation metric $\tilde{I}_{12}$ obtained from purity of subregions for two random sites in the system. (b) Mutual information $I_{12}$ for exact simulations of the mitigated noisy random circuit. (c-d) Scaling collapse of data from (a-b) using $\sigma_{c}/\bar{q}=0.65(5)/0.55(5)$ and critical exponent on $x$-axis obtained from local probe (see Methods A.3). The critical scaling on the $y$-axis is an ansatz. (e-f) Similar quantities as (a-b) for antipodal sites of a 1D chain. The strong drift in the peak with increasing system size is consistent with the lack of a below-threshold phase in this model. In all the plots, we took a Haar random initial pure state with noise parameters $\bar{q}=0.1$ and $p=0.9$ and ran to depth $d=4N$.
II.4 Stability of Ordered Phase to Disorder
@@
To make precise connections of the error mitigation threshold problem to Imry-Ma arguments, we analyze finite-dimensional versions of the statistical mechanics models. First, we define a simple initial condition as a state $\rho_{2}$ that is identity everywhere except in a contiguous region $A$ where it is given by $\operatorname{tr}_{A^{c}}[\rho_{2}]\propto I^{\otimes|A|}+S^{\otimes|A|}$. Such an initial condition can be prepared physically by taking a Haar random state (or a state drawn from a two-design such as the Clifford group DiVincenzo et al. (2 03)) on $A$ and the infinite temperature state on the complement of $A$. The first model we consider is a quenched 1D model described by a one-dimensional brickwork circuit of the type shown in Fig. 1(a) with Haar random two-qubit gates and spatially random noise rates that are constant in time. We define $\sigma=\sqrt{p(1-p)}(q_{2}-q_{1})$ as the standard deviation of the noise rates. We prove that the quenched 1D model always exhibits an instability for one of the simple initial conditions.
@@
II.5 Numerical Tests
Figure 2 summarizes our numerical results. We run numerical simulations of random circuits, together with noise and antinoise, to characterize this phase transition in two settings. In the first case [see Fig. 2(a-d)], we simulate an infinite-dimensional circuit where entangling gates can exist between any pair of qubits and noise is random in both space and time. In the second case [see Fig. 2(e-f)], we simulate $(1+1)D$ circuit dynamics assuming quenched noise, where the randomness is only spatial. The first model is a test of the high-dimensional limit of the model, while the second model is the same one considered in Theorem 1.
In both cases, we initialize a state of a $N$-qubit register in a global Haar-random state, which corresponds to the string $\propto I^{\otimes N}+S^{\otimes N}$ in the statistical-mechanical mapping. For the infinite-dimension model, we apply random Haar gates on a set of $N/2$ disjoint pairs, randomly drawn from all $N(N-1)/2$ possible pairs of qubits. For the $(1+1)D$ model, each time-step corresponds to a layer in the brickwork circuit in Fig. 1 with periodic boundary conditions. Each layer of unitary gates is followed by the noise-antinoise channel on all qubits.
@@ II.5 Numerical Tests
III Discussion
In this work, we showed how imperfect noise characterization in the presence of random spacetime fluctuations in the noise parameters can lead to thresholds in the robustness of error mitigation. We established the absence of a robust phase rigorously in the case of one-dimensional noisy random quantum circuits with quenched spatial disorder. For higher dimensional systems, we presented analytic and numerical evidence that the below threshold phase persists for weak disorder and short enough times. To further support our arguments, we presented a mean-field theory where the phase transition manifests in an instability of the mitigated density matrix to unphysical states.
@@ III Discussion
Regarding asymptotic behavior, the phase transition we discuss only emerges precisely in the thermodynamic limit of large $N$. In this case, a natural concern is that the exponential cost of error mitigation for circuits with depth greater than $\log\log N$ would make the phase transition studied here unphysical Tsubouchi et al. (2023); Takagi et al. (2023); Quek et al. (2022). However, the phase transition studied in this work also arises in constant depth circuits in spatial dimensions of three or greater in the thermodynamic limit. Such behavior was seen directly in the Brownian circuit model analyzed in this work.
An important application of our results is to improved benchmarking and verification of random circuit sampling in noisy devices Boixo et al. (2018); Arute et al. (2019); Bouland et al. (2019); Movassagh (2019); Wu et al. (2021); Zhu et al. (2021). In the Methods A.4, we demonstrate an exponential improvement in a mitigated version of the so-called “linear cross-entropy” fidelity below the error mitigation threshold. Notably, this mitigated benchmark can be computed using existing experimental data. The below threshold behavior may have important consequences for proofs of quantum computational advantage in noisy circuits Aaronson and Gunn (2019); Bouland et al. (2021); Deshpande et al. (2022), potentially motivating the inclusion of error mitigation in complexity theoretic arguments. Moreover, the exponential improvement in the fidelity score for $D\geq 2$ might indicate a method to foil “spoofing” algorithms that have been developed for unmitigated fidelity benchmarks in noisy random circuits Gao et al. (2021).
@@ Acknowledgements.
Code Availability
All code used in this study is available upon request.
@@
All data is available upon request.
Appendix A Methods
A.1 Mean-Field Theory in Brownian Circuits
@@
where $\Delta_{i}=\gamma_{i}-\gamma_{a}$. This equation has the two steady-state solutions $\delta_{i}=0$ and $\delta_{i}=-3/4$, corresponding to the $I^{\otimes N}$ and $S^{\otimes N}$ solutions, respectively. In the case of binary disorder, we define two populations of sites $A_{1/2}$ such that $\Delta_{i}=\gamma_{1/2}-\gamma_{a}$, respectively, for $i\in A_{1/2}$. We have the zero-mean field condition $p\Delta_{1}+(1-p)\Delta_{2}=0$. Defining $G_{\pm}=\frac{1}{N}\Big{(}\sum_{i\in A_{2}}\delta_{i}\pm\sum_{i\in A_{1}}\delta_{i}\Big{)}$, we arrive at simple closed set of equations in the large-$N$ limit
@@
A.2 Statistical Mechanics Mapping Formalism
@@
where $I$ and $S$ are the $4\times 4$ identity matrix and SWAP matrices, respectively. Similarly, the action of a random two-qubit gate on two copies of a qubit-pair is given by
@@
where $I_{2}$ is a $2\times 2$ identity matrix. Similarly,
@@
A.3 Local Probe
A.4 Error Mitigated Fidelity Benchmarks
Here, we introduce mitigated fidelity benchmarks and demonstrate an exponential improvement in a mitigated verison of the linear cross-entropy benchmark below the error mitigation threshold.
$F_{\rm XEB}=2^{n}\sum_{x}p_{n}(x)p_{0}(x)-1,$ (35)
$F_{\rm XEB_{M}}=2^{n}\sum_{x}p_{n}(x)p_{a}(x)-1,$ (36)
$F_{\rm XEB_{M}}=\frac{2^{n}}{M}\sum_{i=1}^{M}p_{a}(x_{i})-1.$ (37)
$\bar{F}_{{\rm XEB}_{M}}=2^{n}\mathbb{E}_{U}\sum_{x}p_{0}(x)p_{an}(x)-1,$ (38)
We can also define a mitigated fidelity that takes the form

### v5

@@ I Introduction
Based on these connections to disordered Ising models, our central prediction is that PEC and TEM have an error mitigation threshold in random circuits in spatial dimensions $D\geq 2$ when the noise is imperfectly characterized, i.e., there is uncertainty in the parameters of the noise model that have random fluctuations in space and/or time. We consider a model of random unitary circuits subject to depolarizing noise with binary disorder in the depolarization rate [see Fig. 1(a)], building on extensive work characterizing quantum phases and phase transitions in hybrid random circuits Fisher et al. (2022); Potter and Vasseur (2022). To mitigate the effects of the noise we apply a uniform “antinoise” channel that inverts the depolarizing noise on average. We show that a replica statistical mechanics description of the problem has close parallels to the classical random field Ising model (RFIM) in $D+1$ dimensions. At the zero-mean field condition in the RFIM, a simple heuristic argument, originally due to Imry and Ma, Imry and Ma (1975) shows that the ordered phase survives random symmetry-breaking terms in sufficiently high dimensions. In the random circuit problem, this analysis indicates the possibility for thresholds in the robustness of error mitigation for $D\geq 2$, while $D=1$ is the marginal dimension. We also analyze the performance of error mitigation as a function of circuit depth. For depths larger than the linear size of the system, the mitigation fails, resulting in a fidelity worse then one obtains for the maximally mixed state. The corresponding conjectured phase diagram of the system is shown in Fig. 1(b) for $D\geq 2$. Analytic and numerical studies of the two-copy replica theory and exact simulations of the mitigated circuit dynamics are consistent with our predictions.
@@
The paper is organized as follows: In Sec. II, we introduce our model for error mitigation based on noise and anti-noise channels. In Sec. III, we present a simple mean-field theory solution for the error mitigation transition in the case of a Brownian circuit model. In Sec. IV, we introduce the statistical mechanics mapping for discrete circuits that we use to analyze the phase transition in finite-dimensional models. In Sec. V, we present a rigorous result on the absence of an ordered phase in 1D with quenched disorder, as well as simple heuristic arguments for the existence of an ordered phase in more than 1D with spacetime random disorder and more than 2D with quenched disorder. In Sec. VI, we present numerical simulations of a mutual information probe in the statistical mechanics model and exactly evolved circuits to support our analytical results. In Sec. VII, we present an alternative probe of the phase transition based on the entropy of a single qubit. In Sec. VIII, we present an application of our results to benchmarking quantum advantage in random circuit sampling experiments. We present our conclusions in Sec. IX. In the appendixes, we present further details on the mean-field calculations and the statistical mechanics mapping.
II Model
@@
[…]
@@
III Mean-field theory
@@
[…]
@@
[…]
@@
where $\Delta_{i}=\gamma_{i}-\gamma_{a}$. This simple equation captures much of the basic physics we find later in the random circuit models, including the presence of a phase transition. Notably, we can see that at weak values of the disorder, the interaction term acts as a restoring force towards the fixed point $\delta_{i}=0$. The other fixed point $\delta_{i}=-3/4$ corresponds to an unphysical density matrix, which is unstable at weak disorder. At a critical disorder strength of $|\Delta_{1}|=3J$ (see App. A), however, the $\delta_{i}=0$ fixed point becomes unstable. The dynamics flows to a new stable fixed point which is an unphysical density matrix. Thus, we see that the above threshold phase in the mean-field theory is characterized by an instabilty of the density matrix to unphysical states.
@@
IV Statistical Mechanics Mapping
To analyze the problem more rigorously, we turn to well-established mappings from random quantum circuit dynamics in $D$ spatial dimensions to statistical mechanics models in $D+1$ dimensions Nahum et al. (2018); von Keyserlingk et al. (2018); Jian et al. (2020); Bao et al. (2020); Dalzell et al. (2022, 2021); Deshpande et al. (2022); Gao et al. (2021). We outline the details of the mapping in Appendix B. The approach we take is to expand $\rho_{k}\equiv\mathbb{U}[\rho^{\otimes k}]$ into a basis of operators based on representations of the permutation group $S_{k}$ Dalzell et al. (2022). One can then derive update rules for the averaged state following each layer of gates, noise, and antinoise. In this paper, we focus on the two-replica case where the two operators are $I$ and $S$ (the SWAP gate). We expect this model to be sufficient to capture many of the properties of the transition, as has been argued in the unitary case Nahum et al. (2018). Numerical evidence based on exact simulations is presented below that support this claim.
@@ IV Statistical Mechanics Mapping
Figure 2: Numerical tests of the threshold behavior: (a) Numerical simulations of circuit and noise-disorder averaged the correlation metric $\tilde{I}_{12}$ obtained from purity of subregions for two random sites in the system. (b) Mutual information $I_{12}$ for exact simulations of the mitigated noisy random circuit. (c-d) Scaling collapse of data from (a-b) using $\sigma_{c}/\bar{q}=0.65(5)/0.55(5)$ and critical exponent on $x$-axis obtained from local probe (see Appendix VII). The critical scaling on the $y$-axis is an ansatz. (e-f) Similar quantities as (a-b) for antipodal sites of a 1D chain. The strong drift in the peak with increasing system size is consistent with the lack of a below-threshold phase in this model. In all the plots, we took a Haar random initial pure state with noise parameters $\bar{q}=0.1$ and $p=0.9$ and ran to depth $d=4N$.
V Stability of Ordered Phase to Disorder
@@
To make precise connections of the error mitigation threshold problem to Imry-Ma arguments, we analyze finite-dimensional versions of the statistical mechanics models. First, we define a simple initial condition as a state $\rho_{2}$ that is identity everywhere except in a contiguous region $A$ where it is given by $\operatorname{tr}_{A^{c}}[\rho_{2}]\propto I^{\otimes|A|}+S^{\otimes|A|}$. Such an initial condition can be prepared physically by taking a Haar random state (or a state drawn from a two-design such as the Clifford group DiVincenzo et al. (2002-03)) on $A$ and the infinite temperature state on the complement of $A$. The first model we consider is a quenched 1D model described by a one-dimensional brickwork circuit of the type shown in Fig. 1(a) with Haar random two-qubit gates and spatially random noise rates that are constant in time. We define $\sigma=\sqrt{p(1-p)}(q_{2}-q_{1})$ as the standard deviation of the noise rates. We prove that the quenched 1D model always exhibits an instability for one of the simple initial conditions.
@@
VI Mutual Information Probe
Figure 2 summarizes our numerical results on the mutual information between two sites in the system, which serves as a probe of the two phases. We run numerical simulations of random circuits, together with noise and antinoise, to characterize this phase transition in two settings. In the first case [see Fig. 2(a-d)], we simulate an infinite-dimensional circuit where entangling gates can exist between any pair of qubits and noise is random in both space and time. In the second case [see Fig. 2(e-f)], we simulate $(1+1)D$ circuit dynamics assuming quenched noise, where the randomness is only spatial. The first model is a test of the high-dimensional limit of the model, while the second model is the same one considered in Theorem 1.
[…]
@@ VI Mutual Information Probe
VII Local Probe
VIII Error Mitigated Fidelity Benchmarks
Here, as an application of our results, we introduce mitigated fidelity benchmarks and demonstrate an exponential improvement in a mitigated verison of the linear cross-entropy benchmark below the error mitigation threshold.
$F_{\rm XEB}=2^{n}\sum_{x}p_{n}(x)p_{0}(x)-1,$ (9)
$F_{\rm XEB_{M}}=2^{n}\sum_{x}p_{n}(x)p_{a}(x)-1,$ (10)
$F_{\rm XEB_{M}}=\frac{2^{n}}{M}\sum_{i=1}^{M}p_{a}(x_{i})-1.$ (11)
$\bar{F}_{{\rm XEB}_{M}}=2^{n}\mathbb{E}_{U}\sum_{x}p_{0}(x)p_{an}(x)-1,$ (12)
We can also define a mitigated fidelity that takes the form
IX Conclusions
In this work, we showed how imperfect noise characterization in the presence of random spacetime fluctuations in the noise parameters can lead to thresholds in the robustness of error mitigation. We established the absence of a robust phase rigorously in the case of one-dimensional noisy random quantum circuits with quenched spatial disorder. For higher dimensional systems, we presented analytic and numerical evidence that the below threshold phase persists for weak disorder and short enough times. To further support our arguments, we presented a mean-field theory where the phase transition manifests in an instability of the mitigated density matrix to unphysical states. Overall, our numerical and analytical results are consistent with the conjectured phase diagram in Fig. 1(b), but further work remains to fully validate the depth-dependence of the phase boundary due to finite-size effects in the numerics.
@@ IX Conclusions
The phase transition we discuss only emerges precisely in the thermodynamic limit of large $N$. In this case, a natural concern is that the exponential cost of error mitigation for circuits with depth greater than $\log\log N$ would make the phase transition studied here unphysical Tsubouchi et al. (2023); Takagi et al. (2023); Quek et al. (2022). However, in more than two dimensions, this phase transition also arises in constant depth circuits, as we showed directly in the case of Brownian circuits. In that case, the phase transition manifests as the emergence of an instability of the mitigated density matrix to unphysical states at large disorder.
An important application of our results is to improved benchmarking and verification of random circuit sampling in noisy devices Boixo et al. (2018); Arute et al. (2019); Bouland et al. (2019); Movassagh (2019); Wu et al. (2021); Zhu et al. (2021). In Sec. VIII, we demonstrated an exponential improvement in a mitigated version of the so-called “linear cross-entropy” fidelity below the error mitigation threshold. Notably, this mitigated benchmark can be computed using existing experimental data. The below threshold behavior may have important consequences for proofs of quantum computational advantage in noisy circuits Aaronson and Gunn (2019); Bouland et al. (2021); Deshpande et al. (2022), potentially motivating the inclusion of error mitigation in complexity theoretic arguments. Moreover, the exponential improvement in the fidelity score for $D\geq 2$ might indicate a method to foil “spoofing” algorithms that have been developed for unmitigated fidelity benchmarks in noisy random circuits Gao et al. (2021).
@@ Acknowledgements.
The data that support the findings of this article are openly available dat .
Appendix A Mean-Field Theory in Brownian Circuits
@@
[…]
@@
Appendix B Statistical Mechanics Mapping Formalism
@@
where $I$ and $S$ are the $\times 4$ 4 identity matrix and SWAP matrices, respectively. Similarly, the action of a random two-qubit gate on two copies of a qubit-pair is given by
@@
where $I_{2}$ is a $\times 2$ 2 identity matrix. Similarly,